# ByteStack-ID: Integrated Stacked Model Leveraging Payload Byte Frequency for Grayscale Image-based Network Intrusion Detection

### v1

@@ I Introduction
The 21st century has ushered in an era of unprecedented era of technological advancement, reshaping our lives through rapid progress in computation and communication technologies [1]. However, this rapid advancement has also expanded the cyber-threat landscape, emphasizing the paramount importance of security in our increasingly interconnected digital world [2]. Furthermore, technologies like the Internet of Things (IoT), where a myriad of low-cost, resource-constrained devices converge on the global stage, further exemplify this trend with an estimated 24.1 billion connected devices projected by 2030, compared to 500 million in 2003 [3]. IoT’s transformative power is evident as it revolutionizes how people and devices interact within connected ecosystems [4].
In this interconnected society, communication and information systems underpin critical infrastructure and personal devices. IoT’s extensive deployment, encompassing nearly 30% of attacks in communication networks [5], poses significant security challenges due to diverse architectures and communication channels [2]. Network Intrusion Detection Systems (NIDS) play a crucial role in identifying and mitigating cyberattacks within a layered defense strategy. Recent advancements emphasize automated cyberattack identification, with machine learning (ML) algorithms gaining traction [6, 7]. In ML, ensemble methods are prominent [8], harnessing the collective power of multiple models like Bagging, Boosting, and Stacking to enhance accuracy [9].
@@ II Related Work
While extensive research has been conducted in the domain of NIDS, a predominant focus has been on flow-based network traffic analysis. While some approaches have delved into packet-level data, addressing the challenge of handling the extensive feature space inherent in packet payloads, we have narrowed our review to such studies that specifically leverage packet-level information.
@@ III Methodology
In this section, we offer a comprehensive overview of our methodology, starting with the dataset we used in our experiments. We then explain the preprocessing steps we took to prepare the data for training and testing, including the conversion of the data into grayscale images. Afterward, we describe our integrated stacked model and how we trained it.
@@ III-A Dataset and Preprocessing
For our study, we utilized the CIC-IDS2017 dataset [18], which was developed by the University of New Brunswick in 2017. This dataset comprises simulated packet-based and bidirectional flow-based network traffic, encompassing modern attacks and benign traffic. The dataset is available in two formats: the original packet capture (PCAP) files (packet-based data) and CSV files (flow-based data). Since we focused primarily on packet-based network traffic, we opted to work with the raw packets from CIC-IDS2017.
@@ III-C1 Architecture of Base Learner
Block#1 BatchNormalization()
Activation(’Relu’)
@@
BatchNormalization()
Activation(’Relu’)
@@
Block#2 BatchNormalization()
Activation(’Relu’)
@@
BatchNormalization()
Activation(’Relu’)
@@
Block#3 BatchNormalization()
Activation(’Relu’)
@@
BatchNormalization()
Activation(’Relu’)
@@ III-C2 Architecture of Meta Learner
For a visual representation of the Integrated Meta Learner, please see Fig. 5. This diagram provides a clear illustration of the integration process within our approach. Furthermore, the comprehensive training process of our approach is outlined in Algorithm 1.
@@ IV Results and Discussion
To ensure a fair comparison, we identified works operating under similar experimental conditions. Detailed results are available in Fig. 6. These results unequivocally demonstrate our approach consistently outperforming other state-of-the-art methods utilizing network traffic packet data. This compelling evidence underscores our approach’s effectiveness and superiority in classifying attack classes using packet level network traffic data.
@@ V Conclusion
Our empirical findings demonstrate ByteStack-ID’s exceptional effectiveness, surpassing baseline models and state-of-the-art approaches. As network threats continue to evolve, ByteStack-ID emerges as a valuable asset in enhancing security measures and proactively mitigating emerging threats within critical network infrastructures. With its pioneering techniques and outstanding performance, ByteStack-ID takes a prominent position at the forefront of intrusion detection research at packet-level network traffic.

### v3

@@ I Introduction
The 21st century has ushered in an era of unprecedented era of technological advancement, reshaping our lives through rapid progress in computation and communication technologies [1]. However, this rapid advancement has also expanded the cyber-threat landscape, emphasizing the paramount importance of security in our increasingly interconnected digital world [2, 3]. Furthermore, technologies like the Internet of Things (IoT), where a myriad of low-cost, resource-constrained devices converge on the global stage, further exemplify this trend with an estimated 24.1 billion connected devices projected by 2030, compared to 500 million in 2003 [4]. IoT’s transformative power is evident as it revolutionizes how people and devices interact within connected ecosystems [5].
In this interconnected society, communication and information systems underpin critical infrastructure [6] and personal devices. IoT’s extensive deployment, encompassing nearly 30% of attacks in communication networks [7], poses significant security challenges due to diverse architectures and communication channels [2]. Network Intrusion Detection Systems (NIDS) play a crucial role in identifying and mitigating cyberattacks within a layered defense strategy. Recent advancements emphasize automated cyberattack identification, with machine learning (ML) algorithms gaining traction [8, 9]. In ML, ensemble methods are prominent [10, 11], harnessing the collective power of multiple models like Bagging, Boosting, and Stacking to enhance accuracy [12].
@@ II Related Work
While extensive research has been conducted in the domain of NIDS, a predominant focus has been on flow-based network traffic analysis [14]. While some approaches have delved into packet-level data, addressing the challenge of handling the extensive feature space inherent in packet payloads, we have narrowed our review to such studies that specifically leverage packet-level information.
@@ III Methodology
This section provides an overview of our methodology, covering the dataset used, preprocessing steps including converting data into grayscale images, and the training of our integrated stacked model, detailing its architecture and workflow.
@@ III-A Dataset and Preprocessing
For our study, we utilized the CIC-IDS2017 dataset [22], which comprises simulated packet-based and bidirectional flow-based network traffic, encompassing modern attacks and benign traffic. The dataset is available in two formats: the original packet capture (PCAP) files (packet-based data) and CSV files (flow-based data). Since we focused primarily on packet-based network traffic, we opted to work with the raw packets from CIC-IDS2017.
@@ III-C1 Architecture of Base Learner
Block#1 BatchNormalization(), Activation(’Relu’)
@@
BatchNormalization(), Activation(’Relu’)
@@
Block#2 BatchNormalization(), Activation(’Relu’)
@@
BatchNormalization(),Activation(’Relu’)
@@
Block#3 BatchNormalization(),Activation(’Relu’)
@@
BatchNormalization(), Activation(’Relu’)
@@ III-C2 Architecture of Meta Learner
For a visual representation of the Integrated Meta Learner, please refer to Fig. 5. This diagram provides a clear illustration of the integration process within our approach. Furthermore, the comprehensive training process of our approach is outlined in Algorithm 1.
@@ IV Results and Discussion
To ensure a fair comparison, we identified works under similar experimental conditions.Detailed results in Fig. 6 unequivocally show our approach consistently outperforming other state-of-the-art methods with network traffic packet data. This underscores the effectiveness and superiority of our approach in classifying attack classes using packet-level network traffic data.
@@ V Conclusion
Our empirical findings demonstrate ByteStack-ID’s exceptional effectiveness, surpassing both baseline models and state-of-the-art approaches. As network threats evolve, ByteStack-ID proves invaluable in enhancing security measures and proactively mitigating emerging threats within critical network infrastructures. With pioneering techniques and outstanding performance, ByteStack-ID leads intrusion detection research in packet-level network traffic.